Maggie0803:

What are the solutions to the equation? 3x^2+6x=45 Enter your answers in the boxes. x = ____ and ____

jhonyy9:

like a first step 3x^2 +(2*3)x = 9*5 so what you can factorizing out from there ?

Maggie0803:

Cancle out the 3?

jhonyy9:

yes divide both sides by 3 and will get what ?

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 3(x^2 +2x) = 3*3*5 what can you doing with both sides now ? \(\color{#0cbb34}{\text{End of Quote}}\)

Maggie0803:

45?

jhonyy9:

no than divide both sides by 3 what will get ? 3(x^2 +2x) = 3*3*5

Maggie0803:

I got it thx

jhonyy9:

np